OpenStudy (anonymous):

Solve the system x - y = 3 and 3x - 8y = 9. What is the value of y? A. 3 B. 0 C. 6 D. -2

Directrix (directrix):

It may be 0 which is option B but I need to check the equation work.

Directrix (directrix):

y = 0

Directrix (directrix):

x - y = 3-----> -3x + 3y = -9 3x - 8y = 9 --->3x -8y = 9 ==================== -5y = 0 y = 0

OpenStudy (anonymous):

1st eqn: y = -3 + x Plug this in to the second eqn: 3x - 8*(-3+x) = 9 3x +24 - 8x = 9 -5x +24 = 9 -5x = -15 x = 3 Plug x=3 into the first eqn: (3) - y = 3 -y = 0 y = 0